# Achieving computational gains with quantum error correction primitives: Generation of long-range entanglement enhanced by error detection

### v1

@@ III.1 State-preparation quality measures
We employ a commonly used characterization method, the multiple-quantum coherence (MQC) Wei et al. (2020); Mooney et al. (2021); Bao et al. (2024), to validate the quality of the GHZ states generated in the experiments described below. We also consider the Hellinger fidelity as a complementary quality measure which is lighter-weight in implementation.
MQC exploits the structure of the GHZ state to measure the fidelity with a number of experiments that scales linearly in system size $n$. Importantly, to establish whether a state exhibits genuine multipartite entanglement, it is sufficient to show $F>0.5$ Leibfried et al. (2005). Here we briefly describe the two quantities measured in MQC: the coherence $C$, and the population $P$. Once obtained, the MQC fidelity is the arithmetic mean of the coherence $C$ and population $P$. Further details on MQC are provided in App. C.
The coherence $C\in[0,1]$ verifies the off-diagonals of the density matrix of a prepared GHZ state. The main idea is to measure the interference pattern (parity oscillation) of a phase-rotated GHZ state, i.e., (|${\left(\ket{00...0}+e^{in\phi}\ket{11...1}\right)/\sqrt{2}}$. Experimentally one measures the interference pattern between the prepared non-phase-rotated and phase-rotated states, $\rho$ and $\rho_{\phi}$, respectively, defined as ${S_{\phi}:=\text{Tr}(\rho\,\rho_{\phi})}$. This quantity should ideally exhibit a sinusoidal behavior as a function of $\phi$ (see Fig. 4(b) for the corresponding measurement circuit). Specifically, the coherence $C$ is related to the $n$-th real Fourier mode, $I_{n}$, of $S_{\phi}$ as
@@ III.2 Experimental results
Next, we disambiguate the role of error suppression and error detection in the experimental results presented above, giving insights into the relative strength of coherent and incoherent errors. For that purpose, we compare the performance of generating a 45Q GHZ state where the same sets of pre-transpiled circuits are executed back-to-back using either error detection and error suppression in combination or error detection alone. The MQC fidelity, Hellinger fidelity, and discarded fraction of shots as a function of number of flags used in post-selection on ibm_brisbane are shown in Fig 5. Both MQC and Hellinger fidelities indicate that when exclusively using error detection (i.e., omitting the error suppression pipeline in circuit execution), GHZ state preparation is dramatically degraded. For example, even when all $5$ flags are used, the MQC fidelity achieved using error detection alone is well below the $0.5$ threshold.
@@ IV Discussion
In this work, we have demonstrated that combining error suppression with QEC primitives can deliver substantial benefits in the computational capabilities of contemporary quantum computers. Through a series of experiments on commercial superconducting processors, we demonstrate record-setting capabilities in both long-range CNOT gates and the generation of maximally entangled GHZ states.
@@
Incorporating the use of QEC primitives delivers a new level of flexibility in the design of circuit modules which exhibit favorable properties for execution on current and near-term devices. For instance, compared to a previously demonstrated measurement-based protocol, the long-range CNOT-gate protocol we introduce trades $n-3$ measurements for the same number of CNOT gates, see Tab. 1. The measurement-based protocol is highly sensitive to single-shot readout errors, since all of the GHZ qubits are in superposition prior to measurement, and the measurement result will inform the feed-forward operations necessary to implement the correct quantum logic. Most contemporary quantum computers exhibit single-shot readout errors that are significantly higher than two-qubit gate errors (by 36×3-6$\times$3 - 6 ×); consequently, our new protocol is better suited for near-term systems. The trade-off comes with a circuit depth that scales linearly with the teleportation path length—this drawback is not expected to be significant for paths with $\lesssim 50$ qubits on current-generation devices, but can considerably limit the gate fidelity with longer teleportation paths (see App. A.2 for scaling the protocol). Moreover, the overall measurement statistics of our protocol will be concentrated on a small number of bit-strings (independent of whether the disentangled qubits are measured), making the measurement distribution amenable to readout error mitigation and providing a further benefit in implementation. This contrasts the inherent ineffectiveness of readout error mitigation for the measurement-based protocol; see App. D.
@@
$n$ ($n$ odd)
@@
Table 1: Overhead for long-range CNOT. Here, $n$ is the size of the shortest-path, linear-chain GHZ state, as well as the distance (measured in qubits) between the control and target qubits of the long-range CNOT. Depth refers to the two-qubit gate depth of the circuit. $N_{\text{CNOT}}$ and $N_{\text{msmt}}$ stand for the number of local CNOT gates and measurements, respectively. Note that we do not count the optional measurements for additional error detection.
@@
Our results make clear that the combination of deterministic error suppression and QEC primitives that stop short of full encoding can provide substantial computational enhancement with far lower overhead than full fault-tolerant QEC. For instance, recent work published while completing this manuscript Reichardt et al. (2024) showed the use of logical encoding of an error-detection code in order to produce a 24-qubit error-detected GHZ state with approximately 90% fidelity; doing so required 59 physical qubits including 35 redundant qubits used in encoding, an overall (including ancillas) qubit overhead ratio of $(59-24)/24\approx 146\%$ and 93% discard rate. In our work a comparable 27-qubit GHZ state also achieving 90% fidelity requires an overhead ratio of only 7% and just 20% shots discarded. At the maximum scale our work uses 84 physical qubits to encode a 75-qubit error-detected GHZ state, requiring only 9 flag ancillas, amounting to a substantially lower overall qubit overhead ratio of $\approx 12\%$ with 78% of shots discarded.
@@
These new results address the fact that most users of contemporary quantum computers will not yet see net computational advantage when leveraging full logical encoding, despite the resounding success of many recent proof-of-principle experiments and the greater generality of logical encoding for arbitrary applications in the long term. Resource overhead in both QPU runtime and hardware redundancy associated with QEC remain challenging and quickly overwhelm the physical device sizes available today. We therefore believe that our observations may help craft guidelines for the design of near-term algorithmic subroutines which deliver expanded computational capability over the best currently achievable routines by adopting the most impactful and efficient QEC primitives.
@@
Looking forward, we suggest two areas for future investigations to further improve the long-range CNOT gate. First, the generation of the linear-chain GHZ can be parallelized via the fusion of smaller GHZ states into a large one connecting the control and target qubits. Second, the disentangling of the GHZ into a distant Bell pair can also be parallelized with multiple branches of unitary disentangling, followed by projective disentangling of the corresponding root qubits. These mixed styles of unitary and projective preparation and disentangling of the linear-chain GHZ allow for a progression from the unitary entangle-disentangle limit ($\mathcal{O}(n)$-depth, $\mathcal{O}(1)$-measurements without error detection) to the measurement-based limit ($\mathcal{O}(1)$-depth, $\mathcal{O}(n)$-measurements): see App. A.2. Such a protocol may yield the optimal balance between two-qubit-gate and readout errors, while restricting the circuit depth, and thus optimizing both the fidelity and the execution speed of the constructed long-range controlled-unitary gate to be applied in e.g., QEC. It also enables the protocol to scale to much longer gate teleportation paths. Finally, we also look forward to leveraging ongoing enhancements being made available by quantum hardware providers, including enhanced support for dynamic circuits IBM (2024).
@@ A.1 Fully unitary disentangling step for long-range CNOT
We recall from the unitary entangle-disentangle protocol described in Sec. II and depicted in Fig. 2 that the intermediate GHZ data qubits are disentangled unitarily following a circuit that mirrors the entangling one—the direction of the local CNOT gates are all towards the ends of the gate teleportation path, leaving the residual Bell pair on the ends untouched. The overall circuit thus resembles the MQC coherence circuit. A consequence of this symmetrical disentangling step is a “root” GHZ data qubit that remains in isolation, which should be projectively disentangled by an $X$-basis measurement. As will become clear below, this disentangling approach is important if one wishes to parallelize the disentangling step in the pursuit of a shallower circuit; see App. A.2.
@@
Figure 7: Parallelization of the entangling and disentangling steps to scale up the long-range CNOT and to optimize for general error budgets and gate speed. (a) An example of fusing a 3-qubit GHZ state (Q0–Q2) and 2 Bell pairs (Q4, Q5, and Q7, Q8) into a 7-qubit GHZ states. The small arrow indicates the direction of the flow of the classical bit. (b) Protocol to create a distant Bell pair over qubits Q0 and Q99. Select qubits are omitted for convenience (indicated by vertical double-arrowed lines and vertically-oriented three open circles), as are certain operations (indicated by horizontal dashed lines). The middle box symbolizes the decoding operation that accepts all the mid-circuit measurement outcomes, and the feedforward operations necessary for the creation of the intermediate GHZ state on all other qubits. At the end of the circuit, the flow of the classical bits of the final measurement outcomes are omitted, since they are only relevant for the parity of the residual Bell pair on Q0 and Q99.
@@ A.2 Designing circuits for general error budgets and scaling up
In summary, a progression from linear-depth, unitary entangle-disentangle protocol, to a constant-depth, measurement-based protocol, while keeping the overall circuit reasonably shallow ($\lesssim 50$ depth), can be realized by parallelizing both the intermediate GHZ state generation and its reduction to a distant Bell pair. The extent of parallelization (namely, the size of smaller GHZ states to be fused, and the size of the subgroup of entangled qubits to be unitarily disentangled) determines the percentage of qubits being measured, allowing for a flexible balance between readout errors and gate errors.
@@ B.1 Average gate fidelity
Figure 12: Adding type-1 parity check flags to the unitary entangle-disentangle long-range CNOT protocol. (a) An example of the long-range CNOT between the control and target qubits (red) across a linear-chain of qubits forming an intermediary GHZ state (green). Type-1 parity-check flags (purple) are inserted to detect errors in the intermediary GHZ state. For example, the parity between Q41 and Q42 is checked by the type-1 flag qubit Q40. (b–d) The long-range CNOT protocol generates a distant Bell pair, benchmarked in terms of (b) fidelity, (c) purity, and (d) concurrence of the reconstructed Bell pair density matrix using post-processed data. Error bars show ±$\pm 1$ standard deviation of $50$ bootstraps.
@@ Appendix C GHZ state preparation with error detection
We then briefly review the MQC formalism. Our aim is to measure the fidelity of the experimentally-prepared and noisy GHZ state, $\tilde{\rho}$. By exploiting the simple structure of the ideal GHZ state, $\rho$, and making an assumption about the form of the noisy state, the MQC formalism allows the fidelity to be estimated using only $32n+$ 3 experiments.
@@ Appendix D Insufficient bit-string counts ineffectuate readout error mitigation
Readout error mitigation techniques that rely on statistical inference Mundada et al. (2023); Nation et al. (2021) are generally ineffective whenever the probability of the most probable bit-string(s) (in the measurement basis), $P_{\text{max}}$, satisfies $P_{\text{max}}\ll 1/N_{\text{shots}}$. For instance, this is satisfied if the underlying (ideal) $n$-qubit state before simultaneous measurements of sufficiently many qubits is a uniform superposition across the exponential number of computational basis states, but is only sampled with a polynomial (in $n$) number of shots. Relevant to our work, three algorithms (circuits) fall into this category and thereby exhibit the phenomenon:

### v2

@@ III.1 State-preparation quality measures
We employ a commonly used characterization method, the multiple-quantum coherence (MQC) Wei et al. (2020); Mooney et al. (2021); Bao et al. (2024), to validate the quality of the GHZ states generated in the experiments described below. We also consider the Hellinger fidelity as a complementary quality measure which is lighter weight in implementation.
MQC exploits the structure of the GHZ state to measure the fidelity with a number of experiments that scales linearly in system size $n$. Importantly, to establish whether a state exhibits genuine multipartite entanglement, it is sufficient to show MQC fidelity $F>0.5$ Leibfried et al. (2005). Here we briefly describe the two quantities measured in MQC: the coherence $C$, and the population $P$. Once obtained, the MQC fidelity $F$ is the arithmetic mean of the coherence $C$ and population $P$. Further details on MQC are provided in App. C.
The coherence $C\in[0,1]$ verifies the off diagonals of the density matrix of a prepared GHZ state. The main idea is to measure the interference pattern (parity oscillation) of a phase-rotated GHZ state, i.e., (|${\left(\ket{00...0}+e^{in\phi}\ket{11...1}\right)/\sqrt{2}}$. Experimentally one measures the interference pattern between the prepared non-phase-rotated and phase-rotated states, $\rho$ and $\rho_{\phi}$, respectively, defined as ${S_{\phi}:=\text{Tr}(\rho\,\rho_{\phi})}$. This quantity should ideally exhibit a sinusoidal behavior as a function of $\phi$ (see Fig. 4(b) for the corresponding measurement circuit). Specifically, the coherence $C$ is related to the $n$-th real Fourier mode, $I_{n}$, of $S_{\phi}$ as
@@ III.2 Experimental results
Next, we disambiguate the role of error suppression and error detection in the experimental results presented above, giving insights into the relative strength of coherent and incoherent errors. For that purpose, we compare the performance of generating a 45Q GHZ state where the same sets of pre-transpiled circuits are executed back-to-back using either error detection and error suppression in combination or error detection alone. The MQC fidelity, Hellinger fidelity, and discarded fraction of shots as a function of number of flags used in post-selection on ibm_brisbane are shown in Fig 5. Both MQC and Hellinger fidelities indicate that when exclusively using error detection (i.e., omitting the error suppression pipeline in circuit execution), GHZ state preparation is dramatically degraded. For example, even when all $5$ flags are used, the MQC fidelity achieved using error detection alone is well below the $0.5$ threshold. We remark that the device ibm_brisbane used here for Fig. 5 has a lower quality compared to the device ibm_fez used for Fig. 4.
@@ IV Discussion
In this work, we have demonstrated that combining error suppression with QEC primitives on unencoded qubits can deliver substantial benefits in the computational capabilities of contemporary quantum computers. Through a series of experiments on commercial superconducting processors, we demonstrate record-setting capabilities in both long-range CNOT gates and the generation of maximally entangled GHZ states.
@@
Incorporating the use of QEC primitives delivers a new level of flexibility in the design of circuit modules which exhibit favorable properties for execution on current and near-term devices. For instance, compared to a previously demonstrated measurement-based protocol, the long-range CNOT-gate protocol we introduce trades $n-3$ measurements for the same number of CNOT gates, see Tab. 1. The measurement-based protocol is highly sensitive to single-shot readout errors, since all of the GHZ qubits are in superposition prior to measurement, and the measurement result will inform the feed-forward operations necessary to implement the correct quantum logic. Most contemporary quantum computers exhibit single-shot readout errors that are significantly higher than two-qubit-gate errors (by 36×3-6$\times$3 - 6 ×); consequently, our new protocol is better suited for near-term systems. The trade-off comes with a circuit depth that scales linearly with the teleportation path length—this drawback is not expected to be significant for paths with $\lesssim 50$ qubits on current-generation devices, but can considerably limit the gate fidelity with longer teleportation paths (see App. A.2 for scaling the protocol). Moreover, the overall measurement statistics of our protocol will be concentrated on a small number of bit-strings (independent of whether the disentangled qubits are measured), making the measurement distribution amenable to readout error mitigation and providing a further benefit in implementation. This contrasts the inherent ineffectiveness of readout error mitigation for the measurement-based protocol; see App. D.
@@
without error detection $n$ ($n$ odd)
Unitary entangle-distentangle $n-1$ ($n$ even) $22n-$ 2 $n$
with error detection $n$ ($n$ odd)
@@
Table 1: Overhead for long-range CNOT. Here, $n$ is the size of the shortest-path, linear-chain GHZ state, as well as the distance (measured in qubits) between the control and target qubits of the long-range CNOT. Depth refers to the two-qubit-gate depth of the circuit. $N_{\text{CNOT}}$ and $N_{\text{msmt}}$ stand for the number of local CNOT gates and measurements, respectively.
@@
Our results make clear that the combination of deterministic error suppression and QEC primitives that stop short of full encoding can provide substantial computational enhancement with far lower overhead than full fault-tolerant QEC. For instance, recent work published while completing this paper Reichardt et al. (2024b) showed the use of logical encoding of an error-detection code in order to produce a 24-qubit error-detected GHZ state with approximately 90% fidelity; doing so required 59 physical qubits including 35 redundant qubits used in encoding, an overall (including ancillas) qubit overhead ratio of $(59-24)/24\approx 146\%$ and 93% discard rate. In our work, although not protected against incoherent phase errors, a comparable 27-qubit GHZ state also achieving 90% fidelity requires an overhead ratio of only 7% and just 20% shots discarded. At the maximum scale, our work uses 84 physical qubits to encode a 75-qubit error-detected GHZ state, requiring only 9 flag ancillas, amounting to a substantially lower overall qubit overhead ratio of $\approx 12\%$ with 78% of shots discarded. Although both kinds of error detection with post-selection have fundamental limits concerning scalability, our adoption of QEC primitives on the physical level has a much lower overall overhead.
@@
These new results address the fact that most users of contemporary quantum computers will not yet see the net computational advantage when leveraging full logical encoding, despite the resounding success of many recent proof-of-principle experiments and the greater generality of logical encoding for arbitrary applications in the long term. Resource overhead in both QPU runtime and qubit count associated with QEC remain challenging and quickly overwhelm the physical device sizes available today. We therefore believe that our observations may help craft guidelines for the design of near-term algorithmic subroutines which deliver expanded computational capability over the best currently achievable routines by adopting the most impactful and efficient QEC primitives.
@@
Looking forward, we suggest two areas for future investigations to further improve the long-range CNOT gate. First, the generation of the linear-chain GHZ can be parallelized via the fusion of smaller GHZ states into a large one connecting the control and target qubits. Second, the disentangling of the GHZ into a distant Bell pair can also be parallelized with multiple branches of unitary disentangling, followed by projective disentangling of the corresponding root qubits. These mixed styles of unitary and projective preparation and disentangling of the linear-chain GHZ allow for a progression from the unitary entangle-disentangle limit ($\mathcal{O}(n)$-depth, $\mathcal{O}(1)$-measurements without error detection) to the measurement-based limit ($\mathcal{O}(1)$-depth, $\mathcal{O}(n)$-measurements): see App. A.2. Such a protocol may yield the optimal balance between two-qubit-gate and readout errors, while restricting the circuit depth, and thus optimizing both the fidelity and the execution speed of the constructed long-range controlled-unitary gate to be applied in e.g., QEC Berthusen et al. (2025). It also enables the protocol to scale to much longer gate teleportation paths. Finally, we emphasize that our novel unitary entangle-disentangle protocol can also serve as a useful QEC primitive itself for bi-layer architectures of solid-state platforms Berthusen et al. (2025).
@@ A.1 Fully unitary disentangling step for long-range CNOT
We recall from the unitary entangle-disentangle protocol described in Sec. II and depicted in Fig. 2 that the intermediate qubits in the GHZ state are disentangled unitarily following a circuit that mirrors the entangling one—the direction of the local CNOT gates are all towards the ends of the gate teleportation path, leaving the residual Bell pair on the ends untouched. The overall circuit thus resembles the MQC coherence circuit. A consequence of this symmetrical disentangling step is a “root” GHZ data qubit that remains in isolation, which should be projectively disentangled by an $X$-basis measurement. As will become clear below, this disentangling approach is important if one wishes to parallelize the disentangling step in the pursuit of a shallower circuit; see App. A.2.
@@
Figure 7: Parallelization of the entangling and disentangling steps to scale up the long-range CNOT and to optimize for general error budgets and gate speed. (a) An example of fusing a 3-qubit GHZ state (Q0–Q2) and 2 Bell pairs (Q4, Q5, and Q7, Q8) into a 7-qubit GHZ states. The small arrow indicates the direction of the flow of the classical bit. (b) Protocol to create a distant Bell pair over qubits Q0 and Q99. Select qubits are omitted for convenience (indicated by vertical double-arrowed lines and vertically oriented three open circles), as are certain operations (indicated by horizontal dashed lines). The middle box symbolizes the decoding operation that accepts all the mid-circuit measurement outcomes, and the feedforward operations necessary for the creation of the intermediate GHZ state on all other qubits. At the end of the circuit, the flow of the classical bits of the final measurement outcomes are omitted, since they are only relevant for the parity of the residual Bell pair on Q0 and Q99.
@@ A.2 Designing circuits for general error budgets and scaling up
In summary, a progression from linear-depth, unitary entangle-disentangle protocol, to a constant-depth, measurement-based protocol, while keeping the overall circuit reasonably shallow ($\lesssim 50$ depth in the near term), can be realized by parallelizing both the intermediate GHZ state generation and its reduction to a distant Bell pair. The extent of parallelization (namely, the size of smaller GHZ states to be fused, and the size of the subgroup of entangled qubits to be unitarily disentangled) determines the percentage of qubits being measured, allowing for a flexible balance between readout errors and gate errors. We finally remark that the advantage of having such a parallelization is absent in SWAP-based methods (e.g., Fig. 6.Ia in Ref. Bäumer et al. (2024)).
@@ A.4 Alternative derivation of the measurement-based protocol for long-range CNOT
Figure 12: Adding type-1 parity check flags to the unitary entangle-disentangle long-range CNOT protocol. (a) An example of the long-range CNOT between the control and target qubits (red) across a linear chain of qubits forming an intermediary GHZ state (green). Type-1 parity-check flags (purple) are inserted to detect errors in the intermediary GHZ state. For example, the parity between Q41 and Q42 is checked by the type-1 flag qubit Q40. (b–d) The long-range CNOT protocol generates a distant Bell pair, benchmarked in terms of (b) fidelity, (c) purity, and (d) concurrence of the reconstructed Bell pair density matrix using post-processed data. Error bars show ±$\pm 1$ standard deviation of $50$ bootstraps.
@@ Appendix C GHZ state preparation with error detection
We then briefly review the MQC formalism. Our aim is to measure the fidelity of the experimentally prepared and noisy GHZ state, $\tilde{\rho}$. By exploiting the simple structure of the ideal GHZ state, $\rho$, and making an assumption about the form of the noisy state, the MQC formalism allows the fidelity to be estimated using only $32n+$ 3 experiments.
@@ Appendix D Insufficient bit-string counts ineffectuate readout error mitigation
Readout error-mitigation techniques that rely on statistical inference Mundada et al. (2023); Nation et al. (2021) are generally ineffective whenever the probability of the most probable bit-string(s) (in the measurement basis), $P_{\text{max}}$, satisfies $P_{\text{max}}\ll 1/N_{\text{shots}}$. For instance, this is satisfied if the underlying (ideal) $n$-qubit state before simultaneous measurements of sufficiently many qubits is a uniform superposition across the exponential number of computational basis states, but is only sampled with a polynomial (in $n$) number of shots. Relevant to our work, three algorithms (circuits) fall into this category and thereby exhibit the phenomenon: